hello mam, A charge Q is divided into two parts of qand Q-q if the coulomb repulsion between them when they are seprated is to be maximum the ratio of Q/q should be
? in this answer its given that for maximum forcedF/dq=0.. may i know how this is taken?

Grade:12

1 Answers

Beenu Mathew
7 years ago
Given a separation r, the force is 

F = k q (Q-q) / r2

This is maximum when q (Q-q) has a maximum. 

The quadratic -q2 + q Q has a maximum where q = -Q/(-2), so when q/Q = ½ or Q/q=2